In a two-digit number, the tens digit is 3 less than twice the ones digit. The number is 27 more than the number obtained by reversing the digits. What is the given number?
Answer by edjones(8007)

Let t=the tens digit. u=the units digit
A) t=2u-3
B) t*10+u=u*10+t+27
.
B) 10(2u-3)+u=10u+2u-3+27 replace t with 2u-3
20u-30+u=12u+24
21u-30=12u+24
9u=54 add 30-12u to each side
u=6 divide each side by 9
.
A) t=2*6-3
t=9
.
the number is 96
.
Check:
96-69=27